OpenStudy (anonymous):

a baseball player hits a baseball into the air with an initial veri]tical velocity of 48 feet per second from a height of 3 feet. write an equation that gives the baseballs height as a function of the time in seconds after it is hit. after how many seconds doesthe baseball reach its maximum height? what is the maximum height? PLEASE HELP

OpenStudy (anonymous):

h(t) = ut - 16t^2 + 3 in your case h(t) = 48t - 16t^2 + 3

OpenStudy (anonymous):

you can find time when height is a maximum by differentiating - have you studied calculus?

OpenStudy (anonymous):

no i havent - algebra:)

OpenStudy (anonymous):

velocity at time t = 48 - 32t at maximum height velocity = 0 so 0 = 48 - 32t t = 1.5 seconds at maxm height

OpenStudy (anonymous):

now to find the maximum height you plug in t = 1,5 in h(t) = 48t - 16t^2 + 3 can you continue with that?

OpenStudy (anonymous):

thnx so much-yes

OpenStudy (anonymous):

you're very welcome